Cashews sell for $4/lb and peanuts sell for $2/lb. How many lbs of each would you use to make 10lbs of a mixture that sells for $2.90/lb?
Answer by jorel1380(3719)

10x2.90=29
p+c=10
2p+4c=29
2p+2c=20
2c=9
c=9/2
p=11/2
You need 11/2 lbs. of peanuts and 9/2 lbs. of cashews..